OpenStudy (anonymous):

Find the integral of sqrt(x^2-2x+1) dx from 0 to 1. (Answer: 1/2)

OpenStudy (turingtest):

factor the integrand

OpenStudy (anonymous):

sqrt(x^2-2x+1)= sqrt(x-1)^2 so that leaves (x-1)dx from 0 to 1

OpenStudy (anonymous):

And then?

OpenStudy (anonymous):

= [(x^2)/2] from 0 to 1 = 1/2

OpenStudy (anonymous):

do you get it?

OpenStudy (anonymous):

So the integral of (x-1) dx is not x^2/2-x?

OpenStudy (anonymous):

sorry im drunk i will tell you now the correct one :D

OpenStudy (anonymous):

its absolute(x-1) because we too sqrt off so from 0 to 1 absoulute(x-1) = -x+1 so the integral becomes [x-(x^2)/2] from 0 to 1 = 1/2 :D

OpenStudy (anonymous):

took*

OpenStudy (anonymous):